OpenStudy (anonymous):

I just have a quick question. What is the difference between a rational and irrational number?

OpenStudy (alexandervonhumboldt2):

rational number is any number that can be expressed as the quotient or fraction p/q of two integers, p and q, with the denominator q not equal to zero.

OpenStudy (alexandervonhumboldt2):

irrational number cannot be expreesed as the quotient or fraction p/q of two integers, p and q, with the denominator q not equal to zero.

OpenStudy (anonymous):

ok thank you

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (perl):

just want to add a note, a number that may not look rational may be rational in fact. For example .11111... is a rational number, because .1111... = 1/9

OpenStudy (perl):

It turns out that decimals which repeat are rational numbers like .13131313... But decimals which don't repeat are irrational, like .121221222... also \(\large \bf \pi \) has a non repeating decimal.

OpenStudy (perl):

but as long as you can take your number and express it as a fraction of two integers, it is a rational number. For example $$ \Large \sqrt{4} =2 = \frac{2}{1} $$ Therefore square root of 4 is rational. Can we do the same for square root of 2?